### v2

@@ A.2 Proof of Proposition 1
A.3 Proof that Partial Orders Dominate Nonpartial Orders
$\varphi((p^{\prime},P^{-\omega_{0}}),\bm{\mathrm{S}})[\omega_{0}]~{}\bar{Q}(\omega_{0})~{}\varphi((p,P^{-\omega_{0}}),\bm{\mathrm{S}})[\omega_{0}]$ for every $P^{-\omega_{0}}$. In addition, we say $\varphi$ is strategy proof with unrestricted lists if submitting the list of acceptable schools that reflects the true preferences weakly dominates submitting anything else for every individual $\omega$.
Lemma 2.
Proof of Lemma 2:
This proof is attributable to haeringer2008, Lemma 4.2. We expand it here in terms of our framework and definitions of partial order.
Second, let $p^{\prime}$ be a weak partial order of $Q(\omega)$ that contains the schools listed in $\bar{p}$. Construct a new “true” preference ranking $q\subseteq\mathcal{J}^{0}$ as follows: (a) take the acceptable schools of $p^{\prime}$, and place them first in $q$ in the same order as they appear in $p^{\prime}$; (b) add a 00 to $q$ after the last school in part (a); (c) fill the remaining positions below 00 in $q$ with the schools not listed in $p^{\prime}$, in any order. Note that $p^{\prime}$ equals the true list of acceptable schools from $q$, $p^{\prime}$ is a weak partial order of $q$, and $p^{\prime}$ is the unique strong partial order of $q$.
The second claim of the lemma follows from the strategyproofness of $\varphi$ since the quota constraint is not binding for an individual whose number of acceptable schools is less than or equal to $K$ and submission of the true list of acceptable schools is feasible.
$\square$
A.4 Proof of Proposition 2
The true preference list $Q$ is unobserved. The submitted preference list $P$ and a weak partial order of $Q$ are observed. Note that all the schools listed in $P$ appear in $Q$ ranked before 00 (i.e., as acceptable schools). There may be other elements in $Q$ not listed in $P$. These remaining schools might appear anywhere in $Q$ as long as the relative ordering of schools in $P$ is preserved in $Q$. Our focus is on students with $P_{j}=(a,b)$, so we know that $a\bar{Q}b$, where $aQb$ if $a\neq b$.
We consider all possibilities of $Q$ that are consistent with the observed $P$ and assumptions and that affect $Q_{j}$. In such cases, the acceptable schools in $Q$ include all the schools in $P$ and possibly more; in fact, since $|P|\leq K<J$, the $J-|P|>0$ unlisted schools in $P$ may appear as acceptable in $Q$. The additional acceptable schools in $Q$ may be schools that are feasible or infeasible within the budget set of the individual.
@@ A.4 Proof of Proposition 2
Regarding the outside option 00, the only way that $Q_{j}$ will have a zero is if $P_{j}$ has a zero. In fact, if $a\neq 0$ and $b\neq 0$, we have that $a\bar{P}bP0$, which implies $a\bar{Q}bQ0$, so none of the coordinates of $Q_{j}$ will be zero. This is why the sets of unlisted feasible options, $N_{j}^{+}$ and $N_{j}^{-}$, do not contain zero.
@@
Consider there is an option $d\in N_{j}^{+}\setminus N_{j}^{-}$. We have that $d\neq a$, $d\neq b$, $d$ is in $B_{j}^{+}(\bm{\mathrm{S}})$ but not in $B_{j}^{-}(\bm{\mathrm{S}})$.
@@
A.5 Proof of Proposition 3
Before the matching mechanism is run, the agent knows her placement scores $\bm{\mathrm{s}}=(s_{1},\ldots,s_{J})$ and her true preferences $Q$ but does not know what the admission cutoffs will be after the matching is run. She sees the admission cutoffs as random variables $(C_{1},\ldots,C_{J})$. The strict preference relation $Q$ is represented by a vector of distinct utility values $(U_{0},U_{1},...,U_{J})$ so that $aQb~{}\Leftrightarrow~{}U_{a}>U_{b}$ for any $a,b\in\mathcal{J}^{0}$. We normalize $U_{0}=0$ for simplicity.
The agent has to decide on a ranking of acceptable schools $P$ to submit. The number of schools ranked in $P$ is $|P|$, and a feasible ranking has $1\leq|P|\leq K$. The set of all feasible rankings is defined as $\Delta P$. We let $P^{u}$ denote the $u$-th school listed in $P$, for $u=1,\ldots,|P|$. We define $L^{P}_{u}$ as the agent’s expected probability of being assigned to school $u$ when submitting ranking $P$, $u=1,\ldots,|P|$. $L^{P}_{0}$ denotes the expected probability of her remaining unassigned, that is, of her taking the outside option. Naturally, $L^{P}_{u}\geq 0$ for every $u$, and $\sum_{u=0}^{|P|}L^{P}_{u}=1$. Cutoff characterization wrt $P$ (Assumption 4) implies:
$\max\limits_{P\in\Delta P}\sum^{|P|}_{u=1}U_{P^{u}}L^{P}_{u}.$
From now on, let $P$ be the optimal choice of this agent. Suppose there is a pair of schools $(d,e)\in UMAS\cap\left(P\times P^{c}\right)$. Let $n$ be the position in $P$ where $d$ appears, i.e., $P^{n}=d$. Construct $\widetilde{P}$ by taking $P$ and replacing option $d$ with option $e$. This implies that $\widetilde{P}^{u}=P^{u}$ for every $u\neq n$ and $\widetilde{P}^{n}=e$.
@@ A.5 Proof of Proposition 3
$L^{P}_{u}=L^{\widetilde{P}}_{u}\text{ for any }u\text{ such that}1\leq u<n\text{ if }n>1,$ (18)
$L^{P}_{n}\leq L^{\widetilde{P}}_{n},$ (19)
$L^{P}_{u}\geq L^{\widetilde{P}}_{u}\text{ for any }u\text{ such that }n<u\leq|P|\text{ if }n<|P|,$ (20)
$L^{P}_{0}\geq L^{\widetilde{P}}_{0},$ (21)
where at least one of the inequalities (19)–(21) is strict if $n<|P|$; or, if $n=|P|$ and at least one of the inequalities (19) and (21) is strict. Below, we prove (18)–(21).
Equation 18 comes from the fact that $\widetilde{P}^{u}=P^{u}$ for every $1\leq u<n$ if $n>1$. The admission probability depends on $u$ being the lowest ranked school for which the agent qualifies. Thus,
@@
For Equation 19, we have that $\mathbb{P}[s_{d}\geq C_{d}]\leq\mathbb{P}[s_{e}\geq C_{e}]$ is implied by the first condition of Assumption 5. This further implies that
@@
For Equation 20, note that the first condition of Assumption 5 implies $\mathbb{P}[s_{d}<C_{d}]\geq\mathbb{P}[s_{e}<C_{e}]$. This further implies that
@@
$\mathbb{P}\left[\cap_{v=1}^{n-1}\{s_{P^{v}}<C_{P^{v}}\}~{}\cap~{}\{s_{e}<C_{e}\}~{}\cap_{v=n+1}^{u-1}\{s_{P^{v}}<C_{P^{v}}\}~{}\cap~{}\{s_{P^{u}}\geq C_{P^{u}}\}\right]=L^{\widetilde{P}}_{u}\par$
@@
For Equation 21, again we have that $\mathbb{P}[s_{d}<C_{d}]\geq\mathbb{P}[s_{e}<C_{e}]$ implies
@@
Finally, the second condition in Assumption 5 (relevance condition) implies that at least one of the inequalities (19)–(21) is strict if $n<|P|$ or, if $n=|P|$, at least one of the inequalities (19) and (21) is strict. Having established these facts, we now move on to compare the expected utility of submitting $P$ to the expected utility of submitting $\widetilde{P}$.
Define $\epsilon=U_{e}-U_{d}=U_{\widetilde{P}^{n}}-U_{P^{n}}>0$. Note that $L_{0}^{P}=1-\sum_{u=1}^{|P|}L^{P}_{u}$ and $L_{0}^{P}-L_{0}^{\widetilde{P}}$ $=\sum_{u=1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)$ $=\left(L^{\widetilde{P}}_{n}-L^{P}_{n}\right)+\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)$, where we adopt the convention that a sum over an empty set of indices equals zero, i.e., $\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)=0$ if $n+1>|P|$. This leads to $\left(L^{\widetilde{P}}_{n}-L^{P}_{n}\right)=\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)-\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)$. Next, we combine these definitions with the inequalities in Equations 1921 to evaluate the difference between the expected utility of submitting $P$ and the expected utility of submitting $\widetilde{P}$:
$\sum^{|P|}_{u=1}U_{P^{u}}L^{P}_{u}-U_{\widetilde{P}^{u}}L^{\widetilde{P}}_{u}=U_{P^{n}}L^{P}_{n}-U_{\widetilde{P}^{n}}L^{\widetilde{P}}_{n}+\sum^{|P|}_{u=n+1}U_{P^{u}}(L^{P}_{u}-L^{\widetilde{P}}_{u})$
$=U_{P^{n}}(L^{P}_{n}-L^{\widetilde{P}}_{n})-\epsilon L^{\widetilde{P}}_{n}+\sum^{|P|}_{u=n+1}U_{P^{u}}(L^{P}_{u}-L^{\widetilde{P}}_{u})$
$=U_{P^{n}}\left[-\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)+\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)\right]-\epsilon L^{\widetilde{P}}_{n}+\sum^{|P|}_{u=n+1}U_{P^{u}}(L^{P}_{u}-L^{\widetilde{P}}_{u})$
$=-U_{P^{n}}\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)-U_{P^{n}}\sum_{u=n+1}^{|P|}\left(L^{P}_{u}-L^{\widetilde{P}}_{u}\right)-\epsilon L^{\widetilde{P}}_{n}+\sum^{|P|}_{u=n+1}U_{P^{u}}\left(L^{P}_{u}-L^{\widetilde{P}}_{u}\right)$
@@
where we use the fact that weak partial order (Assumption 3) implies $U_{P^{1}}>\ldots>U_{P^{|P|}}>0$, $\epsilon>0$, and at least one of $\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)$, $L^{\widetilde{P}}_{n}$, and $\left(L^{P}_{u}-L^{\widetilde{P}}_{u}\right)$ for $u>n$ is strictly positive if $n<|P|$. If $n=|P|$, at least one of $\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)$ and $L^{\widetilde{P}}_{n}$ is strictly positive. The inequality above shows that submitting $\widetilde{P}$ increases the expected utility relative to that from submitting $P$, which contradicts $P$ being the optimal choice. Therefore, $dQe$.
@@ A.6 Proof of Corollary 1
Consider there is an option $f\in N_{j}^{+}\setminus N_{j}^{-}$. We have that $f\neq a$, $f\neq b$, $f$ is in $B_{j}^{+}(\bm{\mathrm{S}})$ but not in $B_{j}^{-}(\bm{\mathrm{S}})$.
@@
$a\bar{P}d$ implies that $a\bar{Q}d$. Given that $dQe$, we have that $a{Q}e$. If $f=e$, we cannot construct $Q$ such that $fQa$. Thus, it is no longer true that we can find $Q$ such that $fQa$ for every $f\in N_{j}^{+}\setminus N_{j}^{-}$; this is true only for $f\neq e$. Therefore, $Q_{j}=(e,b)$ does not belong to $\bm{\mathrm{Q}}_{j}$ in this case.
Case 2.2(b): Suppose $d{P}a$.
$dPa$ implies $dQa$. The fact that $dQe$ does not restrict us from having two possibilities: $dQeQa$ or $dQaQe$. It is again possible to construct $Q$ such that $fQa$ for every $f\in N_{j}^{+}\setminus N_{j}^{-}$, including $f=e$ if $e\in N_{j}^{+}\setminus N_{j}^{-}$. Therefore, $Q_{j}=(e,b)$ does belong to $\bm{\mathrm{Q}}_{j}$ in this case.
@@
A.7 Proof of Proposition 4
The random set $\bm{\mathrm{Q}}_{j}$ is a measurable map from $\Omega$ to $\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 by molinari2020, we say that $\bm{\mathrm{Q}}_{j}$ is a random closed set because for every compact set $\mathcal{K}\in\mathbb{R}^{2}$, the set $\{\omega\in\Omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $Q_{j}$ and the random set $\bm{\mathrm{Q}}_{j}$ are measurable maps on the same probability space, and $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 in molinari2020) characterizes all possible probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$ over $(a,b)\in\mathcal{J}^{0}\times\mathcal{J}^{0}$ for any $Q_{j}$ such that $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$. Since $\bm{\mathrm{Q}}_{j}$ is sharp (Assumption 6(ii)), Artstein’s inequality yields the sharp set of all probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$. For any $s\in\mathcal{S}_{j}$, the inequality says that
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=s\right]\leq\mathbb{P}\left[Q_{j}\in A|S_{j}=s\right]\;\;\;\forall A\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}},$ (22)
@@ A.7 Proof of Proposition 4
Lemma 1 of chesher_ecta2017 applied to our case shows that inequalities (22) are equivalent to:
@@
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~{}|~{}S_{j}=c_{j}^{+}\right]\leq\mathbb{P}\left[Q_{j}\in A~{}|~{}S_{j}=c_{j}\right]\;\;\;\forall A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j},$ (23)
where we use the continuity of $\mathbb{P}\left[Q_{j}\in A~{}|~{}S_{j}=s\right]$ wrt $s$ (Assumption 2) and the existence of the side limit
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~{}|~{}S_{j}=c_{j}^{+}\right]$ (Assumption 7(ii)). Applying an analogous argument to the left of the cutoff $c_{j}$ leads to
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~{}|~{}S_{j}=c_{j}^{-}\right]\leq\mathbb{P}\left[Q_{j}\in A~{}|~{}S_{j}=c_{j}\right]\;\;\;\forall A\in\bm{\mathrm{\Lambda}}^{\cup-}_{j}.$ (24)
If there is $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cap\bm{\mathrm{\Lambda}}^{\cup-}_{j}$, then both (23) and (24) are true, which leads to
$\max\left\{~{}\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{+}\right]~{};~{}\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{-}\right]~{}\right\}\leq\mathbb{P}\left[Q_{j}\in A~{}|~{}S_{j}=c_{j}\right].$ (25)
In summary, we have an inequality for every $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cup\bm{\mathrm{\Lambda}}^{\cup-}_{j}$. There are three possibilities: $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cap\bm{\mathrm{\Lambda}}^{\cup-}_{j}$ (Inequality 25), $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\setminus\bm{\mathrm{\Lambda}}^{\cup-}_{j}$ (Inequality 23), and $A\in\bm{\mathrm{\Lambda}}^{\cup-}_{j}\setminus\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (Inequality 24).

### v3

@@ B.2 Proof of Proposition 1
B.3 Proof of Proposition 2
The true preference list $Q$ is unobserved. The submitted preference list $P$ is observed, and $P$ is a weak partial order of $Q$. Note that all the schools listed in $P$ appear in $Q$ ranked before 0 (i.e., as acceptable schools). There may be other elements in $Q$ not listed in $P$. These remaining schools might appear anywhere in $Q$ as long as the relative ordering of schools in $P$ is preserved in $Q$. Our focus is on students with $P_{j}=(a,b)$, so we know that $a\bar{Q}b$, where $aQb$ if $a\neq b$.
We consider all possibilities of $Q$ that are consistent with the observed $P$ and assumptions and that affect $Q_{j}$. In such cases, the acceptable schools in $Q$ include all the schools in $P$ and possibly more; in fact, since $|P|\leq K<J$, the J|P|>0J-|P|>0 unlisted schools in $P$ may appear as acceptable in $Q$. The additional acceptable schools in $Q$ may be schools that are feasible or infeasible within the budget set of the individual.
@@ B.3 Proof of Proposition 2
Regarding the outside option 0, the only way that $Q_{j}$ will have a zero is if $P_{j}$ has a zero. In fact, if $a\neq 0$ and $b\neq 0$, we have that $a\bar{P}bP0$, which implies $a\bar{Q}bQ0$, so none of the coordinates of $Q_{j}$ will be zero. This is why the sets of unlisted feasible options, $N_{j}^{+}$ and $N_{j}^{-}$, do not contain zero.
@@
Consider there is an option $d\in N_{j}^{+}\setminus N_{j}^{-}$. We have that $d\neq a$, $\neq b, dd$ is in $B_{j}^{+}(\bm{\mathrm{S}})$ but not in $B_{j}^{-}(\bm{\mathrm{S}})$.
@@
B.4 Proof of Proposition 3
The proof is based on the expected-utility maximization problem of Section A.4 and we refer the reader to that section. From now on, let $P$ be the optimal choice of an arbitrary agent with placement scores $\bm{\mathrm{s}}=(s_{1},\ldots,s_{J})$, true preferences $Q$, utilities $(U_{0},U_{1},...,U_{J})$, where $U_{0}=0$, and cutoff beliefs denoted by the random vector $(C_{1},\ldots,C_{J})$.
Suppose there is a pair of schools $(d,e)\in UMAS\cap\left(P\times P^{c}\right)$. Let $n$ be the position in $P$ where $d$ appears, i.e., $P^{n}=d$. Construct $\widetilde{P}$ by taking $P$ and replacing option $d$ with option $e$. This implies that $\widetilde{P}^{u}=P^{u}$ for every $u\neq n$ and $\widetilde{P}^{n}=e$.
@@ B.4 Proof of Proposition 3
[…]
$L^{P}_{n}\leq L^{\widetilde{P}}_{n}$, (B-2)
[…]
$L^{P}_{0}\geq L^{\widetilde{P}}_{0}$, (B-4)
where at least one of the inequalities (B-2)–(B-4) is strict if $n<|P|$; or, if $n=|P|$, at least one of the inequalities (B-2) and (B-4) is strict. Below, we prove (B-1)–(B-4).
Equation B-1 comes from the fact that $\widetilde{P}^{u}=P^{u}$ for every $1\leq u<n$ if $n>1$. The admission probability depends on $u$ being the lowest ranked school for which the agent qualifies. Thus,
@@
For Equation B-2, we have that $\mathbb{P}[s_{d}\geq C_{d}]\leq\mathbb{P}[s_{e}\geq C_{e}]$ is implied by the first condition of Assumption 5. This further implies that
@@
For Equation B-3, note that the first condition of Assumption 5 implies $\mathbb{P}[s_{d}<C_{d}]\geq\mathbb{P}[s_{e}<C_{e}]$. This further implies that
@@
$\mathbb{P}\left[\cap_{v=1}^{n-1}\{s_{P^{v}}<C_{P^{v}}\}~\cap~\{s_{e}<C_{e}\}~\cap_{v=n+1}^{u-1}\{s_{P^{v}}<C_{P^{v}}\}~\cap~\{s_{P^{u}}\geq C_{P^{u}}\}\right]=L^{\widetilde{P}}_{u}$
@@
For Equation B-4, again we have that $\mathbb{P}[s_{d}<C_{d}]\geq\mathbb{P}[s_{e}<C_{e}]$ implies
@@
Finally, the second condition in Assumption 5 (relevance condition) implies that at least one of the inequalities (B-2)–(B-4) is strict if $n<|P|$ or, if $n=|P|$, at least one of the inequalities (B-2) and (B-4) is strict. Having established these facts, we now move on to compare the expected utility of submitting $P$ to the expected utility of submitting $\widetilde{P}$.
Define $\epsilon=U_{e}-U_{d}=U_{\widetilde{P}^{n}}-U_{P^{n}}>0$. Note that $L_{0}^{P}=1-\sum_{u=1}^{|P|}L^{P}_{u}$ and $L_{0}^{P}-L_{0}^{\widetilde{P}}$ $\sum_{u=1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right) =$$\left(L^{\widetilde{P}}_{n}-L^{P}_{n}\right)+\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)$, where we adopt the convention that a sum over an empty set of indices equals zero, i.e., $\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)=0$ if $n+1>|P|$. This leads to $\left(L^{\widetilde{P}}_{n}-L^{P}_{n}\right)=\left(L_{0}^{P}-L_{0}^{\widetilde{P}}\right)-\sum_{u=n+1}^{|P|}\left(L^{\widetilde{P}}_{u}-L^{P}_{u}\right)$. Next, we combine these definitions with the inequalities in Equations B-2B-4 to evaluate the difference between the expected utility of submitting $P$ and the expected utility of submitting $\widetilde{P}$:
[…]
@@
[…]
@@ B.5 Proof of Corollary 1
Consider there is an option $f\in N_{j}^{+}\setminus N_{j}^{-}$. We have that $f\neq a$, $\neq b, ff$ is in $B_{j}^{+}(\bm{\mathrm{S}})$ but not in $B_{j}^{-}(\bm{\mathrm{S}})$.
@@
$a\bar{P}d$ implies that $a\bar{Q}d$. Given that $dQe$, we have that aQea{Q}e. If $f=e$, we cannot construct $Q$ such that $fQa$. Thus, it is no longer true that we can find $Q$ such that $fQa$ for every $f\in N_{j}^{+}\setminus N_{j}^{-}$; this is true only for $f\neq e$. Therefore, $Q_{j}=(e,b)$ does not belong to $\bm{\mathrm{Q}}_{j}$ in this case.
Case 2.2(b): Suppose dPad{P}a.
dPadPa implies $dQa$. The fact that $dQe$ does not restrict us from having two possibilities: $dQeQa$ or $dQaQe$. It is again possible to construct $Q$ such that $fQa$ for every $f\in N_{j}^{+}\setminus N_{j}^{-}$, including $f=e$ if $e\in N_{j}^{+}\setminus N_{j}^{-}$. Therefore, $Q_{j}=(e,b)$ does belong to $\bm{\mathrm{Q}}_{j}$ in this case.
@@
B.6 Proof of Proposition 4
The random set $\bm{\mathrm{Q}}_{j}$ is a measurable map from $\Omega$ to $\mathcal{J}^{0}\times\mathcal{J}^{0}$, so it is compact valued. Following Definition A.1 by Molinari (2020), we say that $\bm{\mathrm{Q}}_{j}$ is a random closed set because for every compact set $\mathcal{K}\in\mathbb{R}^{2}$, the set $\{\omega\in\Omega:\bm{\mathrm{Q}}_{j}(\omega)\cap\mathcal{K}\neq\emptyset\}$ is a measurable event. By Assumption 6(i), the random vector $Q_{j}$ and the random set $\bm{\mathrm{Q}}_{j}$ are measurable maps on the same probability space, and $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$.
Artstein’s inequality (Theorem A.1 by Molinari (2020)) characterizes all possible probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$ over $(a,b)\in\mathcal{J}^{0}\times\mathcal{J}^{0}$ for any $Q_{j}$ such that $\mathbb{P}\left[Q_{j}\in\bm{\mathrm{Q}}_{j}|S_{j}=s\right]=1$ for any $s\in\mathcal{S}_{j}$. Since $\bm{\mathrm{Q}}_{j}$ is sharp (Assumption 6(ii)), the Artstein’s inequality yields the sharp set of all probability mass functions $\mathbb{P}\left[Q_{j}=(a,b)|S_{j}=s\right]$. For any $s\in\mathcal{S}_{j}$, the inequality says that
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=s\right]\leq\mathbb{P}\left[Q_{j}\in A|S_{j}=s\right]\;\;\;\forall A\in 2^{\mathcal{J}^{0}\times\mathcal{J}^{0}},$ (B-5)
@@ B.6 Proof of Proposition 4
Lemma 1 of Chesher and Rosen (2017) applied to our case shows that the inequalities (B-5) are equivalent to:
@@
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~|~S_{j}=c_{j}^{+}\right]\leq\mathbb{P}\left[Q_{j}\in A~|~S_{j}=c_{j}\right]\;\;\;\forall A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}$, (B-6)
where we use the continuity of $\mathbb{P}\left[Q_{j}\in A~|~S_{j}=s\right]$ wrt $s$ (Assumption 2) and the existence of side limit $\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~|~S_{j}=c_{j}^{+}\right]$ (Assumption 7(ii)). Applying an analogous argument to the left of the cutoff $c_{j}$ leads to
$\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A~|~S_{j}=c_{j}^{-}\right]\leq\mathbb{P}\left[Q_{j}\in A~|~S_{j}=c_{j}\right]\;\;\;\forall A\in\bm{\mathrm{\Lambda}}^{\cup-}_{j}.$ (B-7)
If there is $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cap\bm{\mathrm{\Lambda}}^{\cup-}_{j}$, then both (B-6) and (B-7) are true, which leads to
$\max\left\{~\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{+}\right]~;~\mathbb{P}\left[\bm{\mathrm{Q}}_{j}\subseteq A|S_{j}=c_{j}^{-}\right]~\right\}\leq\mathbb{P}\left[Q_{j}\in A~|~S_{j}=c_{j}\right].$ (B-8)
In summary, we have an inequality for every $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cup\bm{\mathrm{\Lambda}}^{\cup-}_{j}$. There are three possibilities: $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\cap\bm{\mathrm{\Lambda}}^{\cup-}_{j}$ (Inequality B-8), $A\in\bm{\mathrm{\Lambda}}^{\cup+}_{j}\setminus\bm{\mathrm{\Lambda}}^{\cup-}_{j}$ (Inequality B-6), and $A\in\bm{\mathrm{\Lambda}}^{\cup-}_{j}\setminus\bm{\mathrm{\Lambda}}^{\cup+}_{j}$ (Inequality B-7).